OpenStudy (anonymous):

.Solve the following trigonometeric equation on the interval [0,2pi). 3cos x = cos x -1 a) pi/2, 4pi/3 b) pi/2, 2pi/3 c) 2pi/3, 4pi/3 d) 4pi/3, 5pi/3 Solvw thw following trigonometric equation on the interval [0, 2pi) a) 5pi/3, 7pi/3, 2pi/3, 4pi/3 b) pi/3, 4pi/3 c) pi/3, 5pi/3 d) 0, pi/3, pi, 5pi/3

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

start with \[2\cos(x)=-1\] so \[\cos(x)=-\frac{1}{2}\] and solve that !